The Evanec Companys next expected dividend, D1, is $3.10; its growth rate is 6%; and its common stock now sells for $36.00.
Answer #1

a.Cost of retained earnings=(D1/Current price)+Growth rate

=(3.1/36)+0.06

=14.61%(Approx).

b.% flotation cost=(Current price-Net price)/Current price

=(36-34.2)/36

=5%

c.Cost of new common stock=(D1/Net price)+Growth rate

=(3.1/34.2)+0.06

=15.06%(Approx).